# ShanghaiTech Mapping Robot is All You Need: Robot System for Collecting Universal Ground Vehicle Datasets

### v1

@@ I Introduction
  • System paper of very nice SLAM dataset collection robot
  • Calibration
  • Carachterization
  • Dataset with Evaluation
Figure 1: The ShanghaiTech Mapping Robot II.
We are presenting a UGV-based mapping robot capable of accommodating, powering, synchronizing and controlling various kinds of sensors orientated towards multiple directions as well as collecting and storing data from all sensors. The robot is designed to accommodate (i) facing the front direction, a pair of wide-angle color camera, a pair of infra-red camera, a pair of event-based camera, an RGB-Depth camera, a mmWave radar, 2 solid-state non-rotational LiDARs, (ii) facing each of the left, right, back and up direction, a pair of wide-angle color camera, an RGB-Depth camera, a mmWave Radar, (iii) facing downward, one wide-angle color camera, (iv) on the top, an polydioptric omnidirectional color camera (consisting of 6 wide-angle color camera) and 2 real-time kinematic GNSS antennas, (v) surrounding the robot, 5 rotational LiDARs, 13 ultrasound range finder and an AHRS IMU.
The vehicle mechanically consists of the UGV base, structure for sensor mounting, the power box and the computation cluster.
II Related Work
In recent years, several public datasets have been introduced to support research in autonomous driving and mobile robotics. The KITTI dataset [1] is one of the pioneering works that provides synchronized camera and LiDAR data captured from a moving vehicle. The dataset includes RGB images, grayscale stereo images, 3D Velodyne point clouds, and GPS/IMU data.
The Cityscapes dataset [2] focuses on semantic understanding of urban street scenes. It provides stereo RGB images with pixel-level and instance-level semantic annotations for 30 classes. The dataset is captured in 50 different cities during daytime and good weather conditions.
The Oxford RobotCar Dataset [3] introduces a large-scale autonomous driving dataset captured over a year in Oxford, UK. The dataset covers different times of day, weather conditions, and traffic densities. It includes RGB images, 2D LiDARs, 3D LiDARs, GPS/IMU, and RTK ground truth.
More recent datasets such as nuScenes [4], Waymo Open Dataset [5], and Argoverse [6] incorporate richer sensor modalities and provide larger data volumes. The nuScenes dataset includes RGB cameras, LiDARs, radars, and GPS/IMU. Waymo Open Dataset and Argoverse provide high-quality LiDAR point clouds and 3D object annotations.
While these datasets have significantly advanced research in autonomous driving, they are still limited in certain aspects. Most datasets focus on a specific sensor configuration and do not cover the full spectrum of sensing modalities. The calibration quality and synchronization accuracy are often not rigorously evaluated. Most importantly, the diversity of environments and conditions is still limited.
Our ShanghaiTech Mapping Robot seeks to address these limitations by incorporating a more comprehensive sensor suite, ensuring high-quality calibration and synchronization, and capturing data in more diverse environments. The dataset and benchmarks enabled by our robot will complement existing datasets and support new research directions in robot autonomy.
III Fleet of Sensors
@@
Ground vehicles like road automotive, warehouse AGVs, or even household robot vacuum cleaners, all have different sensors equipped by them, suiting their size, environment of operation and budget. Moreover, the choice of sensor is affected by the indented mission of the vehicle. Most vehicles would use the data for basic localization and mapping while some of them also expect fine reconstruction of environment or detecting surrounding objects they try to avoid bumping into (or contrarily, would like to interact with).
@@
Different types of sensors suit different tasks. To achieve the goal of collecting universal ground vehicle datasets, we tried our best to equip our robot with almost all kinds of sensors utilized in SLAM and self-driving field, hoping that the datasets it produce would fit most, if not all, SLAM algorithms. To be specific, our choice covers visual sensors (including frame-based RGB cameras, depth cameras, infra-red cameras and event-based cameras), LiDARs (including rotational ones and solid ones), inertial sensors, mmWave radars, ultrasonic range finders and GNSS receivers. In this section, we will introduce the fleet of sensors on the robot.
@@ III-A RGB Cameras
RGB camera is one of the most used sensor in the SLAM field, providing data for visual SLAM algorithms and occupancy or semantics networks. Conventional algorithms utilize either mono or stereo visual streams. State-of-the-art occupancy networks in auto-driving field even use the bird-eye-view image stitched from surrounding cameras as the only input for the network. Thus we equipped the mapping robot with 2 models of RGB cameras, FLIR Grasshopper3 (GS3-U3-51S5C-C) and FLIR LadyBug5+.
FLIR Grasshopper3 (GS3-U3-51S5C-C) is a 2/3-inch color camera that has a resolution of $2448\times 2048$ pixel and can capture images in global shutter mode at the frame rate up to 75fps.[7] Our initiative was to use them to achieve stereo view towards five main directions: front, back, left, right, up, and to have a monocular downward view of the floor or ground the robot drives on. FLIR LadyBug5+ is a polydioptric omnidirectional color camera, consisting of 6 pinhole cameras. 5 of the cameras face the surrounding while the last one faces upward. Each of the 6 sensors has a resolution of $2464\times 2048$ and a maximum frame rate of 30fps [8].
Figure 2: EvC3A event-based camera, FLIR Grasshopper3 (GS3-U3-51S5C-C), and Azure Kienct DK depth camera (from left to right respectively)
@@ III-B RGB-Depth Cameras
We chose Azure Kinect DK as the RGB-D camera of our robot. It features a 12MP high-resolution ($3072\times 1920$) RGB camera that can capture color images in global shutter mode at 30fps and an 1MP depth camera that uses ToF technology to measure distance within the range of 0.5 - 5.46 meters [9]. Compared to depth cameras that use structured light technology, ToF RGB-D cameras work better in outdoor environments and have a longer range, which is more suitable for our robot.
@@
Figure 3: A315 infra-red camera (left) and Ladybug 5+ polydioptric camera (right)
@@ III-C IR Cameras
For our robot, we chose FLIR A315, a compact uncooled microbolometer thermal camera. A315 features a $320\times 240$ pixels Vanadium Oxide detector, having a temperature detection range from -20°C to 120°C with an accuracy up to ±2°C. Its built-in lens provides a 25° field of view and motorized focusing control. It has a spatial resolution of 1.36 mrad and a temperature resolution of 50 mK. Two A315s should be used to form a forward looking stereo pair.
@@ III-D Event-based Cameras
We equipped the robot with EvC3A event-based cameras featuring the Prophesee PPS3MVCD sensor (often being referred as the Gen3.1 VGA Sensor) also used in the EVK3 camera. It has a resolution of $640\times 480$ on its 3/4” sensor, achieving a typical latency of 200µs and a dynamic range above 120dB. Two of them should be deployed to form a forward looking stereo pair.
@@ III-E LiDARs
In order to diversify the technical methodologies employed by the LiDAR systems and facilitate comprehensive coverage of the sensing area in all directions, we equipped the robot with seven LiDARs, encompassing five distinct models: Robosense Ruby, Robosense Bpearl, Hesai PandarQT, Livox Avia and Neuvition M1.
Figure 4: Robosense Bpearl, Robosense Ruby, and Hesai PandarQT LiDARs (from left to right respectively)
Robosense Ruby and Hesai PandarQT are conventional rotational LiDARs. Robosense Ruby is a 128-beam LiDAR with a 0.1° angular resolution on both the horizontal and vertical axes, 200m range and 40° vertical field of view, well-suited for autonomous driving applications requiring far detection range. In contrary, Hesai PandarQT, also known as QT64, is a 64-beam LiDAR designed for indoor applications. It features a 60m range, 0.6° horizontal angular resolution and has a vertical angular resolution up to 1.45°. With a 104.2° (-52.1° to +52.1°) vertical aperture, its FOV covers a typical room from the ceiling to the floor.
Robosense Bpearl is a 32-beam rotational LiDAR featuring a hemispherical $360^{\circ}\times 90^{\circ}$ field of view. It has a small blind range of less than 0.1m and a maximum range of 30m, designed for near-field detection. It provides a vertical resolution of 2.81° and a horizontal resolution up to 0.2°.
Figure 5: Livox Avia (left) and Neuvition M1 (right) LiDARs
Livox Avia and Neuvition M1 are unconventional LiDARs. Livox Avia is a semi-solid-state LiDAR featuring a unique non-repetitive scanning pattern and a forward-looking field of view. Unlike conventional rotational LiDARs, Livox Avia’s coverage area ratio increases with the time of integration. It has a $70.4^{\circ}\times 77.2^{\circ}$ field of view, 190m range and ±2cm accuracy. It features a best point density among all 5 LiDARs euipped by our robot and thus fits use cases like fine mapping and 3D reconstruction. Neuvition M1 is a MEMS-based solid-state lidar, providing a $45^{\circ}\times 25^{\circ}$ field of view with a $1750\times 480$ resolution and 200m range at 20% reflectivity.
@@ III-F IMUs
We chose MTi-630R, a rugged and compact attitude and heading reference system (AHRS) from Xsens, to be the main IMU of our robot. It provides 0.2° RMS roll/pitch and 1° RMS heading accuracy and is able to perform high precision strapdown integration at 400Hz. It also measures temperature and atmospheric pressure, potentially helps determine the elevation of robot indoor.

### v2

@@ I Introduction
Figure 1: The ShanghaiTech Mapping Robot
Localization is an essential task of all mobile robots - they need to know their pose for path planning and navigation, as well as to perform their actual task. While there exist localization methods that use external infrastructure, such as GPS or special indoor markers [1], the most general and flexible way of localization is to localize in a map [2]. Additionally, maps themselves are also needed for path planning and navigation. Thus, mapping is also a very important task for many mobile robotic systems. Typically, we need to solve those two problems together, because localization needs the map and mapping needs localization, with algorithms called Simultaneous Localization and Mapping (SLAM) [3]. Being such an essential algorithm, there is lots of research on SLAM. For proper scientific evaluation of the performance of SLAM algorithms, authors need to compare their solution to other, state-of-the-art approaches. For that, datasets of recorded robotic sensor data are used and then we typically compare the ground truth robot poses of these datasets with the poses estimated by the SLAM algorithms.
One major problem with this approach is, that there are various SLAM methods that use a multitude of different sensors and combination thereof, e.g. LiDAR, monocular camera, stereo, depth camera, event camera, radar, ultrasound, with or without IMU or odometry data. Those sensors may also be pointed in different directions or feature different resolutions or frame rates. In order to test and compare their solutions, SLAM researchers need to choose datasets compatible with their approach. But that means, that they can only compare their algorithm with other methods that work with similar datasets. Often researchers even need to collect their own dataset, suitable for their SLAM approach. The consequence of all this is, that currently, most SLAM papers only compare to a very limited number of other state-of-the-art algorithms, and it is thus often unclear how well their solution performs compared to many other, state of the art methods.
So ideally, there should be a dataset that can be used by all SLAM algorithms. Such a dataset would feature all kinds of different sensors in many different configurations, with very high specifications (e.g. resolution and frame-rate), such that SLAM algorithms can pick and choose which sensors they want, while potentially sub-sampling the data to fit their need. It is also not sufficient to do recording sessions with the different sensors in the same environment - what is needed for proper fair, repeatable and reproducible evaluation of the algorithms is, that all sensors should follow the exact same trajectory. To collect such a dataset we thus need a robot that can collect all this different sensor data in one go. In this paper we present our efforts to build such a robot: The ShanghaiTech Mapping Robot, which is all the mapping robot needed for ground vehicle dataset collection.
The requirements for this dataset collection system are as follows: It should be a system for indoor and outdoor environments and as compact as possible. Because many sensors should be supported, which are in total quite heavy, flying systems, legged as well as hand-held solutions are impossible, so a wheeled mobile robotics base has to be chosen. While still being reasonable in terms of size, weight and consumption of power and computation resources, the robot should feature 1) as many different types sensors as possible with 2) as high specifications as possible and 3) with as big fields of view as possible. The ShanghaiTech Mapping Robot thus features 36 sensors of 9 different modalities. The amount of data collected by all of these sensors can be up to 3,000 MiB per second. To facilitate the collection and storage of all this data we equipped our robot with a 16-node cluster. The robot is depicted in Fig. 1.
This work was inspired by our first mapping dataset collection robot, which was much simpler and featured far fewer sensors [4]. In [5] we already reported on the design, construction and experiments of the 16-node ”Cluster on Wheels”. The ultimate goal of this work is to collect datasets that can be utilized by the SLAM Hive Benchmarking Suite [6, 7] to properly evaluate SLAM algorithms by performing 10s of thousands mapping runs and analyzing their results in-depth.
@@
  • In-depth discussion of the requirements, design and construction of the ultimate dataset collection robot;
    • Mechanical Design
    • Electrical System
    • Sensor Setup
    • Computing System
    • Distributed Software System
    • Hardware-Synchronization
  • Calibration approach, esp. extrinsic calibration of 32 sensors
  • Sensor Interference Evaluation
  • Outdoor Dataset with SLAM Experiments
The rest of this paper is organized as follows: Section II reviews related work on SLAM datasets and multi-sensor robotic platforms. Section III describes the sensors of our data collection system, while section IV discusses the basic backbone systems of the robot and section V introduces the hardware and software integration of the sensors with the cluster as well as the hardware synchronization system. The intrinsic and extrinsic calibration procedures of the sensors are explained in section VI. Section VII then reports on the data collection and post-processing procedures. A few sensors interfere with each other and the limited effect of this is investigated and described in section VIII. The outdoor dataset provided with the paper, as well as SLAM experiments with it, are presented in section IX. This paper concludes with section X, also discussing the future work. Details of the data collected by the robot are provided in Appendix A.
II Related Works
The development of acquisition of SLAM datasets has been crucial for advancing research in robotics, computer vision, and autonomous driving. This section reviews relevant datasets and techniques, focusing on multi-sensor approaches, visual SLAM, LiDAR-based systems, event-based cameras, and mmWave radar applications.
Several large-scale datasets have been instrumental in driving progress in autonomous vehicle research. The KITTI dataset [8], introduced in 2013, was one of the pioneering efforts, providing synchronized stereo imagery, LiDAR scans, and GPS/IMU data. Building upon this foundation, the Oxford RobotCar dataset [9] offered long-term data collection over varying conditions, while the Cityscapes dataset [10] focused on semantic understanding of urban scenes. More recent contributions like nuScenes [11] and Waymo Open Dataset [12] have further expanded the scope, incorporating additional sensor modalities and more diverse driving scenarios. The Ford Multi-AV Seasonal Dataset [13] specifically addresses the challenges of seasonal variations. Our work distinguishes itself from these datasets by offering a more extensive and diverse sensor suite, including event-based cameras, infra-red cameras, a combination of 5 different LiDARs models, and a multi-directional mmWave radars setup, which are not commonly found in existing datasets.
LiDAR approaches are one of the prevalent SLAM methods, as well as for 3D reconstruction and autonomous driving applications. Datasets such as M2DGR [14] and the Multi-Modal Lidar Dataset [15] have provided valuable resources for developing and evaluating LiDAR-based SLAM algorithms. The RUGD dataset [16] specifically targets unstructured environments, while PandaSet [17] offers a comprehensive sensor suite including LiDAR for autonomous driving research. Our work builds upon these efforts by integrating high-resolution LiDAR data with other sensor modalities, allowing for more robust and versatile perception algorithms.
Visual SLAM, as yet another key approach of SLAM, has seen significant advancements, fundamentally owing to the availability of diverse datasets. The EuRoC MAV datasets [18] provided benchmark data for aerial vehicles, while more recent works like the multi-camera framework by Kaveti et al. [19] and the visual-inertial odometry system by Zhang et al. [20] have explored multi-camera setups for improved robustness. The Campus Map dataset [21] offers a large-scale multi-camera dataset with additional sensors, similar to our approach. However, our work extends beyond this by incorporating event-based cameras and mmWave radars, enabling research into fusion algorithms that can leverage these novel sensing modalities.
Event-based cameras have emerged as a promising technology for high-speed and high-dynamic-range applications. Datasets such as ViViD++ [22], DSEC [23], and VECtor [24] have provided valuable resources for developing event-based SLAM and odometry algorithms. The Event-Camera Dataset and Simulator [25] and the Multivehicle Stereo Event Camera Dataset [26] have further expanded the available data for event-based research. Our work incorporates event-based cameras alongside traditional sensors. It allows for the development of hybrid algorithms like [27] that can exploit the complementary strengths of different sensor modalities while also creates a benchmark for various event-based algorithms to compete with their visual or lidar-based counterparts.
The use of mmWave radar in SLAM and autonomous driving is an emerging field with significant potential. Recent works such as the multi-radar inertial odometry by Huang et al. [28] and the 4D imaging radar inertial odometry and mapping (iRIOM) by Zhuang et al. [29] have demonstrated the capabilities of mmWave radar for ego-motion estimation and mapping. Park et al. [30] proposed a novel radar velocity factor for pose-graph SLAM, while Lu et al. [31] developed milliMap for robust indoor mapping using mmWave radar. Our robot includes mmWave radar data, enabling research into multi-sensor fusion algorithms that can leverage the unique properties of radar, such as detecting glass, using the Doppler effect to meature object’s speed and its robustness in adverse weather conditions.
In summary, while existing datasets have made significant contributions to various aspects of SLAM and autonomous driving research, our work uniquely combines a wide array of sensor modalities, including event-based cameras and mmWave radars. This comprehensive approach allows for the development and evaluation of more robust and versatile algorithms capable of operating in diverse and challenging environments.
III Sensor Suite
@@
Ground vehicles like road automotive, warehouse AGVs, or even household robot vacuum cleaners, all have different sensors equipped, suiting their size, environment of operation and budget. Moreover, the choice of sensor is affected by the indented mission of the vehicle. Most vehicles would use the data for basic localization and mapping, while some of them also expect fine reconstruction of environment or detecting surrounding objects they try to avoid bumping into (or contrarily, would like to interact with).
@@
Different types of sensors suit different tasks. To achieve the goal of collecting universal ground vehicle datasets, we tried our best to equip our robot with almost all kinds of sensors utilized in SLAM and self-driving field, hoping that the datasets it produce would fit most, if not all, SLAM algorithms. To be specific, our choice covers visual sensors (including frame-based RGB cameras, depth cameras, infra-red cameras and event-based cameras), LiDARs (including rotational ones and solid ones), inertial sensors, mmWave radars, ultrasonic range finders and GNSS receivers. In this section, we will introduce the sensor suite on the robot.
Figure 2: Sensors on ShanghaiTech Mapping Robot.
TABLE I: Sensor Configuration of the ShanghaiTech Mapping Robot
Sensor Type Model Qty Specifications Max Data Rate Field of View Looking Direction Appearance
RGB Camera FLIR Grasshopper3 11 2448 x 2048 pixels 75 fps 81.9° x 61.2° Front, Back, Left, Right, Up (stereo); Down (mono) Fig. 2-b
RGB Camera FLIR LadyBug5+ 1 (6) 6 x 2464 x 2048 pixels 30 fps 90% of full sphere Omnidirectional Fig. 2-a
RGB-D Camera Azure Kinect DK 5 RGB: 3072 x 1920, Depth: 1MP 30 fps N/A Front, Back, Left, Right, Up Fig. 2-c
IR Camera FLIR A315 2 320 x 240 pixels 60 Hz 25° Front (stereo pair) Fig. 2-e
Event-based Camera EvC3A 2 640 x 480 pixels N/A N/A Front (stereo pair) Fig. 2-d
Spinning LiDAR Robosense Ruby 2 128 beams, 200m range 10 Hz 360° x 40° Horizontal, Vertical (90° pitched) Fig. 2-i
Spinning LiDAR Hesai PandarQT 1 64 beams, 60m range 10 Hz 360° x 104.2° Horizontal Fig. 2-h
Spinning LiDAR Robosense Bpearl 2 32 beams, 30m range 10 Hz 360° x 90° Front, Back Fig. 2-j
Solid-state LiDAR Livox Avia 1 190m range 10 Hz 70.4° x 77.2° Front Fig. 2-k
Solid-state LiDAR Neuvition M1 1 200m range 10 Hz 45° x 25° Front Fig. 2-l
IMU Xsens MTi-630R 1 400 Hz update rate 400 Hz N/A N/A Fig. 2-g
Ultrasonic Sensor KS103 bundle 1 1cm to 8m range N/A 13 different angles Front Fig. 2-m
mmWave Radar TI IWR6843AOPEVM 5 60GHz N/A 120° x 120° Front, Back, Left, Right, Up Fig. 2-n
GNSS Receiver Bynav A1 1 Dual antenna with IMU N/A N/A N/A Fig. 2-f
Total 36 (41) 17 RGB, 5 RGB-D, 2 Event-based, 2 IR, 7 LiDARs, 5 mmWave Radars, 1 Sonar Set, 1 IMU, 1 GNSS RTK/INS
@@ III-A RGB Cameras
Color cameras are some of the most used sensors in the SLAM field, providing data for visual SLAM algorithms, as well as providing input for advanced computer vision algorithms like semantic segmentation or object detection. Conventional algorithms such as ORB-SLAM3 utilize either monocular or stereo visual streams [32].
Bird’s-Eye-View data, stitched from omni-directional cameras, is also very popular recently, e.g. [33].
Thus we equipped the mapping robot with 2 models of RGB cameras, FLIR Grasshopper3 (GS3-U3-51S5C-C) and FLIR LadyBug5+.
FLIR Grasshopper3 (GS3-U3-51S5C-C) (Fig. 2-b) is a 2/3-inch color camera that has a resolution of $2448\times 2048$ pixel (5MP) and can capture images in global shutter mode at the frame rate up to 75fps.[34] Our initiative was to use them to achieve stereo view towards five main directions: front, back, left, right, up, and to have a monocular downward view of the floor or ground the robot drives on. FLIR LadyBug5+ (Fig. 2-a) is a polydioptric omnidirectional color camera, consisting of 6 pinhole cameras. 5 of the cameras face the surrounding while the last one faces upward. Each of the 6 sensors has a resolution of $2464\times 2048$ and a maximum frame rate of 30fps [35].
@@ III-B RGB-Depth Cameras
We chose Azure Kinect DK (Fig. 2-c) as the RGB-D camera of our robot. It features a 12MP high-resolution ($3072\times 1920$) RGB camera that can capture color images in global shutter mode at 30fps and an 1MP depth camera that uses ToF technology to measure distance within the range of 0.5 - 5.46 meters [36]. Compared to depth cameras that use structured light technology, ToF RGB-D cameras work better in outdoor environments and have a longer range, which is more suitable for our robot.
@@ III-C IR Cameras
For our robot, we chose FLIR A315 (Fig. 2-e) , a compact uncooled microbolometer thermal camera. A315 features a $320\times 240$ pixels Vanadium Oxide detector, having a temperature detection range from -20°C to 120°C with an accuracy up to ±2°C. Its built-in lens provides a 25° field of view and motorized focusing control. It has a spatial resolution of 1.36 mrad and a temperature resolution of 50 mK. Two A315s should be used to form a forward looking stereo pair.
@@ III-D Event-based Cameras
We equipped the robot with EvC3A event-based cameras (Fig. 2-d) featuring the Prophesee PPS3MVCD sensor (often being referred as the Gen3.1 VGA Sensor) also used in the EVK3 camera [37]. It has a resolution of $640\times 480$ on its 3/4” sensor, achieving a typical latency of 200µs and a dynamic range above 120dB. Two of them should be deployed to form a forward looking stereo pair.
@@ III-E LiDARs
In order to diversify the technical methodologies employed by the LiDAR systems and facilitate comprehensive coverage of the sensing area in all directions, we equipped the robot with seven LiDARs, encompassing five distinct models: Robosense Ruby , Robosense Bpearl , Hesai PandarQT , Livox Avia, and Neuvition M1 .
Robosense Ruby and Hesai PandarQT are conventional rotational LiDARs. Robosense Ruby (Fig. 2-i) is a 128-beam LiDAR with a 0.1° angular resolution on both the horizontal and vertical axes, 200m range and 40° vertical field of view, well-suited for autonomous driving applications requiring far detection range. In contrary, Hesai PandarQT (Fig. 2-h) , also known as QT64, is a 64-beam LiDAR designed for indoor applications. It features a 60m range, 0.6° horizontal angular resolution and has a vertical angular resolution up to 1.45°. With a 104.2° (-52.1° to +52.1°) vertical aperture, its FOV covers a typical room from the ceiling to the floor. Both of them, as well as all other LiDARs equipped by our robot, run in dual-return or triple-return mode, which can improve reflection detection accuracy [38].
Robosense Bpearl (Fig. 2-j) is a 32-beam rotational LiDAR featuring a hemispherical $360^{\circ}\times 90^{\circ}$ field of view. It has a small blind range of less than 0.1m and a maximum range of 30m, designed for near-field detection. It provides a vertical resolution of 2.81° and a horizontal resolution up to 0.2°.
Livox Avia and Neuvition M1 are unconventional LiDARs. Livox Avia (Fig. 2-k) is a semi-solid-state LiDAR featuring a unique non-repetitive scanning pattern and a forward-looking field of view. Unlike conventional rotational LiDARs, Livox Avia’s coverage area ratio increases with the time of integration. It has a $70.4^{\circ}\times 77.2^{\circ}$ field of view, 190m range and ±2cm accuracy. It features a best point density among all 5 LiDARs euipped by our robot and thus fits use cases like fine mapping and 3D reconstruction. Neuvition M1 (Fig. 2-l) is a MEMS-based solid-state lidar, providing a $45^{\circ}\times 25^{\circ}$ field of view with a $1750\times 480$ resolution and 200m range at 20% reflectivity.
@@ III-F IMUs
We chose MTi-630R (Fig. 2-g) , a rugged and compact attitude and heading reference system (AHRS) from Xsens, to be the main IMU of our robot. It provides 0.2° RMS roll/pitch and 1° RMS heading accuracy and is able to perform high precision strapdown integration at 400Hz [39]. It also measures temperature and atmospheric pressure, potentially helps determine the elevation of robot indoor.